### v1

# Tuning Event Camera Biases Heuristic for Object Detection Applications in Staring Scenarios

Abstract
One of the main challenges in unlocking the potential of neuromorphic cameras, also called “event cameras”, is the development of novel methods that solve the multi-parameter problem of adjusting their bias parameters to accommodate a desired task. Actually, it is very difficult to find in the literature a systematic heuristic that solves the problem for any desired application.
In this paper we present a tuning parametes heuristic for the biases of event cameras, for tasks that require small objects detection in staring scenarios. The main purpose of the heuristic is to squeeze the camera’s potential, optimize its performance, and expand its detection capabilities as much as possible.
In the presentation, we translate the experimental properties of event camera and systemic constrains into mathematical terms, and show, under certain assumptions, how the multi-variable problem collapses into a two-parameter problem that can be solved experimentally.
A main conclusion that will be demonstrated is that for certain desired signals, such as the one provided by an incandescent lamp powered by the periodic electrical grid, the optimal values of the camera are very far from the default values recommended by the manufacturer.
Key words: neuromorphic camera, event camera, biases tuning, optimization problem, control system, detection application.
AMS subject classification: 49J21, 93C35, 93B52, 93C65.
1 Introduction
Event camera is a bio-inspired sensor that pose a paradigm shift in acquisition of visual information. Event camera gives information about changes in the field of view. Each pixel of an event camera is asynchronously independent and responds only when it feels a brightness change that reaches a predefined threshold. Hence, in contrast to standard cameras that sample the whole field of view in a certain sampling-rate and output a sequence of synchronous frames, event cameras provide data only when a certain pixel feels a predifined change in the illumination, and the output is a list of asynchronous polarized "events" sorted by their timestamp. Event cameras come with some special properties: timestamp in resolution of microseconds, low latency, high dynamic range (over 120 dB), and low power consumption (see [1, 2] for review).
As event cameras work differently from frame-based cameras, a main challenge in unlocking their potential is to develop novel asynchronous event driven methods and algorithms to process their output (see [3, 4, 5, 6, 7, 8, 9, 10, 11] for some examples).
However, in addition to the different output of the camera, the degrees of freedom of event camera are different from those of frame-based camera, and also less intuitive. For example, when using a conventional camera, one needs to suit the frame rate and exposure time to the desired task. When desciding about the frame rate, we need to consider the amount of data that the system can deal with, and when desciding about the exposure time, one needs to think about avoiding glare and blurring that can happen by overexposing from one hand, but also to avoid underexposing as well. However, event camera does not have these two degrees of freedom at all. Instead, it has a few biases that control the generation of the events. Two biases control the sensitivity of the pixels to changes in the illumination. One bias controls an internal low-pass-filter in the pixels, that is usually considered as the bias that is responsible for controling the bandwidth of the camera [1]. Another bias controls an internal high-pass-filter, this bias is usually ignored in the literature. The last bias controls the refractoty period of the pixels. These five degrees of freedom make the tuning procedure into a multi-variable problem that one needs to solve in order to exploit the camera’s potential for any application. In particular, when answering the question whether event camera can give an added value with relation to a frame-based camera, it is crucial to first solve the aforementioned problem and optimize the biases with relation to the desired task. Although one can find technical information and methods on tuning the biases in the literature [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], it is difficult to find an approach that deals with this critical challenge as a multi-variable problem.
In this paper we are going to present a heuristic for tuning the event camera’s biases for small objects detection applications in staring scenarios. By “object detection in staring scenarios” we assume the following assumptions:
  1. 1.
    The camera is staring, i.e. the field of view of the camera does not change during the recording. In particular, the event rate of the camera is stable when the desired signal is missing. When we say “stable” we mean that the event rate might vary from time to time, but not by orders of magnitude.
  2. 2.
    Our goal is to detect the object in the highest significance we can. In other words, we refer to the biases as optimal with relation to the SNR in an appropriate predefined terms.
In the literature, one can find specific discussions about the dependence of the camera performance and noise on the biases [13, 15, 16, 17, 18, 19, 20, 23]. Some of it is devoted to the agreed requirement of achieving similar sensitivity to positive and negative events, and how to get to it under the technical limitations of the pixels operation [17, 19]. We accept this requirement, as discussed below, and we also add the requirement of having constant event rate [14] (see $\oint$3). However, in contrast, we do not focus on the technical details on how to acheive these requirements [12, 14, 21, 22]. We assume that it is possible, even though we are aware to the limitations of this statement. Instead, we focus on the interpretation and meaning of these two requirements regarding the desired mission, and how to step forward to the solution of the tuning biases multi-variable problem.
In the discussion below we describe schematically the influence of the biases on the event rate and interpret these properties into mathematical properties of the multi-variable problem (see $\oint$4). This description is based, inter alia, on a few conclusions from our experience with the Prophesee’s EVK4-HD event camera model. Under this interpretation, we prove the following existence and uniqueness theorem (see $\oint$5):
Theorem 1.
Let $f(x,y)$ and $g(x,y)$ be real valued differentiable functions, defined over the rectangles $[a,b)\times[c,d]$ and $[a,b]\times[c,d)$, correspondingly. Assume also that for every $(x_{0},y_{0})$ in the regions we have:
$\underset{x\to a}{\lim}f(x,y_{0})$ $=\underset{y\to c}{\lim}g(x_{0},y)=0$
$\underset{x\to b}{\lim}f(x,y_{0})$ $=\underset{y\to d}{\lim}g(x_{0},y)=\infty$
$\frac{\partial f(x_{0},y_{0})}{\partial x}$ $>\frac{\partial g(x_{0},y_{0})}{\partial x}>0$
$\frac{\partial g(x_{0},y_{0})}{\partial y}$ $>\frac{\partial f(x_{0},y_{0})}{\partial y}>0$
  1. 1.
    Then, for every constant $0<C\in\mathbb{R}$ there exists exactly one solution to the equations system
    $f(x,y)=g(x,y)=C.$
  2. 2.
    If $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are the unqiue solutions of the system for the constants $0<C_{1}<C_{2}$, correspondingly, then $x_{1}<x_{2}$ and $y_{1}<y_{2}$.
In the sequel, when we use the notion of “event camera” we actually refer to those models that have the degrees of freedom regarding the aforementioned tuning parameters, putting aside event cameras that do not have them. To make the presentation of the heuristic even more clear and concrete, we specify our discussion on EVK4-HD Prophesee’s model. Though, the same heuristic can be applied to other relevant models under certain adjustments.
Using the aforementioned theorem, we show that the multi-variable problem collapses into two-dimensional problem that can be studied experimentally. Then, in $\oint$6 we conclude by showing how one can study the two-dimensional problem for a periodic light signal powered by the electricity grid. A very interesting result of the analysis is that the optimal values for the biases are very different from the default values recommended by Prophesee.
2 The biases of an event camera
Contrary to a frame-based camera, the output of an event camera does not consist of a synchronous series of matrices that describe the gray level of the pixels. Instead, it consists of an asynchronous list of “events” that are generated in the following way.
Any pixel of the camera “remembers” a certian reference value for the intensity of the light in the pixel. Then, the pixel measures continuously the change of the intensity of the light with relation to the reference value. When the intensity of the light changes enough, and the change reaches a predefined value, the pixel takes two actions: It updates the reference value to the current intensity value, and reports an “event”. Each event reported by the camera contains the following information: The timestamp of the change, in resolution of microseconds; The coordinates of the pixel; The polarity of the event, namely, if the event was triggered by a positive change of the light or a negative one. The list of the events, which is the output of the camera, is given to the user, sorted by the timestamps of the events.
Event camera has 5 parameters that control the generation of the events. Each one of the parameters can be tuned upward and downward. Tuning the parameter to one direction causes the camera to report more events regarding the signal, but also to report more events from the background and to increase the noise. Tuning the parameter to the other direction causes the opposite. In the following we present these 5 parameters.
Two parameters are dedicated to control the thresholds for triggering the events - one parameter for triggering positive events and one for negative. In Prophesee’s user interface these parameters are called “bias_diff_on” and “bias_diff_off”. We will denote them by $P$ (for “Positive”) and $N$ (for “Negative”), respectively. Both parameters can be tuned separately, and in general, lowering them causes the thresholds to be smaller, and the camera to be more sensitive, i.e. to collect more data from the signal and to increase the noise.
Another two parameters are dedicated to control internal low pass (LPF) and high pass filters (HPF) of the pixels. Each pixel, by its photoreceptor, collects photons from its field of view, and then translates the optic signal to an electric signal. But before passing the electric signal to the unit that decides whether there was a sufficient change in the illumination to report an event, the pixel passes it through a LPF and a HPF. The parameters that control the frequencies range of the LPF and HPF are denoted in Prophesee’s interface by “bias_fo” and “bias_hpf”, respectively. We will denote them by $L$ (for “LPF”) and $H$ (for “HPF”).
In general, increasing the latter two parameters tunes the camera to pass higher frequencies of the signal and to block lower ones. In other words, increasing $L$ means that we allow more high frequencies to pass, causing the camera to report more events from the signal, but to increase background event rate (i.e. the noise) as well. From the other hand, increasing $H$ blocks more small frequencies, causing the camera to report less events from the signal, and to reduce the noise. In general, there is not an obvious known way to connect between these two biases and the signal frequencies blocked by the filters.
The last parameter controls the refractory period of the pixels, which is the time between the pixel reported its last event, and the time that the pixel will be able to report another one. The main purpose of this feature is to avoid the camera from reporting events that are caused by internal noise of the pixel after reporting an event. In addition, it reduces the event rate caused by hot pixels, i.e. sensitive pixels that report relatively a large amount of events that are caused by internal noise. This feature is not intrinsic to the way event camera works, and idealy, this issue should be solved by improving the stabilization of the pixels themselves. Even now, one can see that in EVK4-HD this is almost non-issue, and the default refractory period is quite low, ranging between 10-100 microseconds [18], while the event rate caused by hot pixels is negligible. Therefore, we do not refer to this parameter in this paper.
Hence in general we are left with 4 degrees of freedom in our multivariable optimization problem: two for the parameters that control the sensitivity of the camera, and two for the parameters that control the internal LPF and HPF.
3 Main principles of the heuristic
3.1 Principle 1: fixing the event rate
In applications such as those we are dealing here, one can assume that once the biases were tuned to a certain point, the event rate is stable. That means that when there is no motion in the field of view, the event rate is close to be constant. In this case, the events in the camera are mostly caused by sporadic inner noise of the pixels, and by vibrations of edges in the field ot view, caused by the turbulence of the air. When we use the notion of “background event rate”, or $BER$, we refer to this constant event rate. This is different from the momentary event rate in the camera that is proportional to the motion of objects in the field of view. For example, the momentary event rate can increase substantially if a car passes near the camera and covers its whole field of view in a certain moment.
In general, the main problem in squeezing the biases so that the camera will be as sensitive as possible, and pass as much frequencies from the signal as possible, is increasing the background event rate of the camera, and the momentary event rate. Too high event rate can be problemtic in a few aspects:
  1. 1.
    When the event rate is very high, it harms the stability of the event stream, and the data that the camera reports becomes less reliable. High event rate can also influence the real time operation of the camera.
  2. 2.
    When the camera reports a lot of events, it can become very difficult to process the data in terms of real time algorithms and computation resources.
  3. 3.
    When the camera reports a lot of events near the desired signal, it becomes more difficult, algorithmically, to distinguish the signal from the events in the background. In contrast to the two former aformentioned problem, this aspect can cause a classical SNR problem.
Considering the first aspect, in our reference camera, Prophesee’s EVK4-HD, from our experience, it seems that the camera works smoothly up to more or less 1M events/sec, and one should avoid getting into the area of 10M events/sec. Hence, systemically, one should consider the $BER$ to be in the range of up to 100K events/sec, and to be flexible enough and keep the option of the staring event camera to reach 1M events/sec event rate in case of substantial momentary unexpectable motions in the field of view. So, from now on, we assume that systemically, one should not tune the biases to a certain point that allows the fixed background event rate to substantially surpass the value of 100K events/sec.
Moving to the second aspect, after fixing $BER\leq$100K events/sec, from our experience, asynchronous event-driven algorithms, such as in [11], can be easily implemented in real time using available computers. However, this property is not essential to the presented heuristic, as if the implementation of the detection algorithm requires a smaller $BER$ value, denoted by $D$, then one can replace the restriction $BER\leq$100K events/sec by $BER\leq D$. So, without loss of generality, we assume that we have a certain constant event rate value $D$, smaller than (or equal to) 100K events/sec, that we do not want to surpass, from systemic considerations.
Regarding the last aspect, one should notice that when the background event rate is smaller than 100K events/sec, the averaged event rate of a single pixel in an HD-camera is ~0.1 events/sec. Assuming that the amount of events triggered by the desired signal is at least a few, say 5 events, it is quite clear that we do not have here a classical SNR problem. In other words, it might be difficult to distnguish between our desired signal and other real motions in the field of view, but not from the events triggered by the background.
Concluding the discussion up to here, we assume that one should tune the biases so that the background event rate will not surpass the value of $D\leq$100K events/sec. However, as under this constraint, as was explained above, there is no real noise, the optimal SNR will be achieved by optimizing the signal itself. Let us now denote the function that describes the size of the signal by
$Sig(L,H,P,N).$
A general definition for $Sig$ can be the amount of events triggered by the desired object. In the demonstration below we will define it a bit different, and count the amount of positive or negative events, and for some applications it can be something else. The main property of $Sig$ is that we want to optimize its value with relation to $L,H,P,N$ under the constraint $BER\leq D$.
3.2 Principle 2: equalizing positive vs negative sensitivity
In general, depending on the desired signal, one can think of different options for the desired realtion between $N$ and $P$. However, in general, as discussed in [17], it seems that for most applications, one would like that the sensitivity of the camera to positive and negative events will be equal. We say that, even about detection missions in which the desired signal consists of events of single polarity, say positive events. The main reason for that is the connection between the ability of the pixels to report positive events and negtive events, due the way event camera works. Assume, for example, the extreme case in which the desired signal consists of solely positive events. Hence, one can think of tuning the camera to be very sensitive to positive events, and to be non-sensitive at all to negative events. Now, how in general the pixel of event camera works? As was described above, it keeps a certain reference value for the contemporary illumination intensity, and it asks whether the change with relation to that reference value is big enough in order to report an event. Hence, if the camera is not sensitive to negtive events at all, then, given a certain pixel, after a while, the illumination in the pixel might grow as a result of certain irrelevant temporary motions in the field of view, and as a result, the pixel will report some positive events, and will update its reference illumination value to be higher. However, even after these motions are gone, the reference value will not be updated to a lower one as the sensitivity to negative events is very low. Therefore, when the desired signal will come, the reference value will be already high, and even the positive desired events will not be reported as the signal migth not be strong enough to surpass the reference value. So in order for an event camera to react in a stable way for positive or negative events, it should be able to react for both polarities.
Now, one might think that there is a direct way to tune the biases so that the sensitivity for both polarities will be equal, but practically, there is not. It seems that the only available way to balance the camera’s sensitivity for both polarities, is to tune $N$ and $P$ in such a way that the positive $BER$ and negative $BER$ are close to be equal. Practically, they cannot be completely equal, but we are going to show why this requirement does work in general, and usually one can tune the biases and get quite close to that.
So, in conclusion, we agree that in general, the most reasonable heuristic for tuning the sensitivity biases, is to try to have the same sensitivity for negative and positive events. In addition, we agree that the most practical way to do so is to try tuning the sensitivity such that the positive $BER$ and negative $BER$ are equal. Lastly, in the sequel we are not going to present a specific algorithm that equalizes the positive and negative $BER$, we are going to explain why such an algorithm will work. One can find such implementations and discussions on how to do it in [12, 14, 21, 22].
4 Biases tuning heuristic
Given a specific scene, for each value of $(L,H,P,N)$ we will get a certain value for the positive $BER$ and negative $BER$. Let’s denote these values by $BER_{P}(L,H,P,N)$ and $BER_{N}(L,H,P,N)$, respectively. Notice that
$BER(L,H,P,N)=BER_{P}(L,H,P,N)+BER_{N}(L,H,P,N)$
and that one of our principles is that $BER_{P}(L,H,P,N)=BER_{N}(L,H,P,N)$.
Now, as in general the information from the background grows when the camera is more sensitive, one can assume that $BER_{P}$ and $BER_{N}$ are monotonically decreasing as function of $P$ and $N$. In other words for every $L,H,P,N$ one has
$N<N^{\prime}$$\Rightarrow\begin{cases}BER_{N}(L,H,P,N)>BER_{N}(L,H,P,N^{\prime})\\ BER_{P}(L,H,P,N)>BER_{P}(L,H,P,N^{\prime})\end{cases}$
$P<P^{\prime}$$\Rightarrow\begin{cases}BER_{N}(L,H,P,N)>BER_{N}(L,H,P^{\prime},N)\\ BER_{P}(L,H,P,N)>BER_{P}(L,H,P^{\prime},N)\end{cases}$
We remark that at a first glance, it is not trivial that changing the value of $P$ will affect $BER_{N}$ and vice versa. However, as was discussed at the beginning of $\oint$3.2, one can notice both idealy and experimentally, that the change of the sensitivity in one polarity changes the $BER$ of both polarities. However, one can certainly notice that when changing the sensitivity for one polarity affects the $BER$ of the corresponding polarity more that it affects the $BER$ of the other polarity.
Now, for the sake of the discussion, let us assume that $BER_{N}(L,H,P,N)$ and $BER_{P}(L,H,P,N)$ are differentiable with relation to their parameters. In particular, $N$ and $P$ can be tuned continuously. In reality, the values of $N$ and $P$ are discrete, but conceptually, there is no reason to assume it when thinking about the biases multi-variable problem.
Hence, one can assume that for every $L,H,P,N$ we have
$\frac{\partial BER_{N}}{\partial N}$ $<\frac{\partial BER_{P}}{\partial N}<0$
$\frac{\partial BER_{P}}{\partial P}$ $<\frac{\partial BER_{N}}{\partial P}<0$
which encodes the aforementioned property that a change in $P$ affects $BER_{P}$ more than it affects $BER_{N}$, and vice versa for $N$.
Now, let $N_{min},N_{max},P_{min},P_{max}$ be the limits of the parameters $N$ and $P$ in the camera. Then, another idealy right property is that for every $L,H,P,N$ one has
$\underset{n\to N_{max}}{\lim}BER_{N}(L,H,P,n)$ $=\underset{p\to P_{max}}{\lim}BER_{P}(L,H,p,N)=0$
$\underset{n\to N_{min}}{\lim}BER_{N}(L,H,P,n)$ $=\underset{p\to P_{min}}{\lim}BER_{P}(L,H,p,N)=\infty$
which means that no matter what, if we make the camera non sensitive for a certain polarity, then the appropriate $BER$ will vanish, and if we tune the camera to the opposite direction, the appropriate $BER$ will explode.
We remark that in reality, sometimes the first raw is not correct as one can have scenarios with highly active background, and the maximal values for $N$ and $P$ are not sufficient to reduce the $BER$ to 00. However, in most cases it can be achieved, and in the others it can be achieved for a certain region of $L$ and $H$ values.
Now, given a pair $(L,H)$, one can denote
$f(x,y)$ $=BER_{N}(L,H,-x,-y)$
$g(x,y)$ $=BER_{N}(L,H,-x,-y)$
defined over the region $(-N_{max},-N_{min})\times(-P_{max},-P_{min})$. Then, $f$ and $g$ satisfy the conditions of Theorem 1. Hence, for every $C>0$ there exists a unique pair $(X,Y)=(-N,-P)$ such that $f(X,Y)=g(X,Y)=C$, or equivalently
$BER_{N}(L,H,N,P)=BER_{P}(L,H,N,P)=C.$
Now, recall the constraint $BER\leq D$. Recall also that as was explained above, under this constraint we are not dealing with SNR in the usual manner, and our mission is to find the optimal $(L,H,P,N)$ for reaching the highest signal. Hence, as one can assume that the information from the signal grows when the camera is more sensitive, we can assume also here that $Sig$ is monotonically decreasing as function of $P$ and $N$. In other words, for every $L,H,P,N$ one has
$N<N^{\prime}$ $\Rightarrow Sig(L,H,P,N)>Sig(L,H,P,N^{\prime})$
$P<P^{\prime}$ $\Rightarrow Sig(L,H,P,N)>Sig(L,H,P^{\prime},N).$
Corollary 2.
Under the above assumptions, for every pair $(L,H)$, the optimal value of $Sig$, under the constraints
$BER(L,H,N,P)$ $\leq D$
$BER_{N}(L,H,N,P)=BER_{P}(L,H,N,P)$
is given by the unique pair $(N^{\prime},P^{\prime})$ that satisfies
$BER_{N}(L,H,N^{\prime},P^{\prime})=BER_{P}(L,H,N^{\prime},P^{\prime})=D/2.$
Proof.
Assume negatively that $Sig(L,H,N^{\prime\prime},P^{\prime\prime})$ is higher than $Sig(L,H,N^{\prime},P^{\prime})$ and
$BER(L,H,N^{\prime\prime},P^{\prime\prime})$ $\leq D$
$BER_{N}(L,H,N^{\prime\prime},P^{\prime\prime})=BER_{P}(L,H,N^{\prime\prime},P^{\prime\prime})$
By the uniqeness property of $(N^{\prime},P^{\prime})$, it follows that $BER(L,H,N^{\prime\prime},P^{\prime\prime})=D^{\prime}<D$. Hence, by the second part of Theorem 1 we have $N^{\prime}<N^{\prime\prime}$ and $P^{\prime}<P^{\prime\prime}$. It follows that $Sig(L,H,N^{\prime},P^{\prime})>Sig(L,H,N^{\prime\prime},P^{\prime\prime})$, and we have a contradiction. ∎
The above discussion yields that in order to optimize the signal, it is enough to examine the $(L,H)$-plane where $N(L,H)$ and $P(L,H)$ are the unique values of $N$ and $P$ that Theorem 1 and Corollary 2 promise. In other words, all we need to do is to evaluate
$\tilde{Sig}(L,H)=Sig(P(L,H),N(L,H),L,H)$
We are going to demonstrate it in the last section.
Remark 3.
In general, as the pixels of event cameras are logarithmic, the sensitivity of the pixels is directly affected by changes in the illumination. As a result, changes in the illumination, automatically changes the $BER$. Assuming that the filters are much less affected by the illumination change, when staring on a certain view, after one finds the optimal values of $L$ and $H$ for the signal, when the illumination changes, all is needed to be done, is to tune $N$ and $P$ to the new unique value that will lead to satisfy the conditions of Theorem 1 and Corollary 2.
5 Proof of Theorem 1
Following the assumptions
$\underset{y\to c}{\lim}g(x,y)$ $=0$
$\underset{y\to d}{\lim}g(x,y)$ =$\infty$= ∞
and the monotony of $g$ with relation to $y$, by the intermediate value theorem, it follows that for every $a\leq x<b$ there exists exactly one value $y(x)$ such that
$g(x,y(x))=C.$ (5.1)
Now, we want to show that the function $f(x,y(x))$ is monotonically increasing as function of $x$, even though $y$ is not fixed.
Using the formula for the derivative of the implicit function $y(x)$ given by Equation 5.1, one has
$\frac{\partial y}{\partial x}=-\frac{\partial g}{\partial x}/\frac{\partial g}{\partial y}.$
Hence, by the chain rule and the assumptions of the theorem on the derivatives of $f$ and $g$, one has
$\frac{\partial f(x,y(x))}{\partial x}$ $=\frac{\partial f}{\partial x}+\frac{\partial f}{\partial y}\cdot\frac{\partial y}{\partial x}$
$=\frac{\partial f}{\partial x}-\frac{\partial f}{\partial y}\cdot\left(\frac{\partial g}{\partial x}/\frac{\partial g}{\partial y}\right)\geq\frac{\partial f}{\partial x}-\frac{\partial g}{\partial x}>0$
yielding that $f(x,y(x))$ is monotonically increasing as required. Now, notice that
$\underset{x\to a}{\lim}f(x,y(x))$ $\leq\underset{x\to a}{\lim}f(x,d)=0$
$\underset{x\to b}{\lim}f(x,y(x))$ $\geq\underset{x\to b}{\lim}f(x,c)=\infty$
and therefore
$\underset{x\to a}{\lim}f(x,y(x))$ $=0$
$\underset{x\to b}{\lim}f(x,y(x))$ =$\infty$= ∞
Using the intermediate value theorem once again, for the monotonically increasing function $f(x,y(x))$, it follows that there exists exactly one $x$ such that
$f(x,y(x))=C.$
This gives us the first part of Theorem 1. For the second part, let us write the equations
$f(x(C),y(C))=g(x(C),y(C)=C$
where $x(C)$ and $y(C)$ are the unique values given by the first part for each $C$. Then, for the implicit functions $x(C)$ and $y(C)$ we have, by the chain rule
11 11 $=\frac{\partial f}{\partial C}=\frac{\partial f}{\partial y}\cdot\frac{\partial y}{\partial C}+\frac{\partial f}{\partial x}\cdot\frac{\partial x}{\partial C}$
11 11 $=\frac{\partial g}{\partial C}=\frac{\partial g}{\partial y}\cdot\frac{\partial y}{\partial C}+\frac{\partial g}{\partial x}\cdot\frac{\partial x}{\partial C}$
Hence
$\left(\begin{array}[]{cc}\frac{\partial f}{\partial x}&\frac{\partial f}{\partial y}\\ \frac{\partial g}{\partial x}&\frac{\partial g}{\partial y}\end{array}\right)\cdot\left(\begin{array}[]{c}\frac{\partial x}{\partial C}\\ \frac{\partial y}{\partial C}\end{array}\right)=\left(\begin{array}[]{c}1\\ 1\end{array}\right)$
yielding that
$\left(\begin{array}[]{c}\frac{\partial x}{\partial C}\\ \frac{\partial y}{\partial C}\end{array}\right)=\frac{1}{\frac{\partial f}{\partial x}\cdot\frac{\partial g}{\partial y}-\frac{\partial f}{\partial y}\cdot\frac{\partial g}{\partial x}}\cdot\left(\begin{array}[]{cc}\frac{\partial g}{\partial y}&-\frac{\partial f}{\partial y}\\ -\frac{\partial g}{\partial x}&\frac{\partial f}{\partial x}\end{array}\right)\cdot\left(\begin{array}[]{c}1\\ 1\end{array}\right)$
and therefore, by the assumptions $\frac{\partial g}{\partial y}>\frac{\partial f}{\partial y}>0$ and $\frac{\partial f}{\partial x}>\frac{\partial g}{\partial x}>0$ we have
$\frac{\partial x}{\partial C}$ $=\frac{\frac{\partial g}{\partial y}-\frac{\partial f}{\partial y}}{\frac{\partial f}{\partial x}\cdot\frac{\partial g}{\partial y}-\frac{\partial f}{\partial y}\cdot\frac{\partial g}{\partial x}}>0$
$\frac{\partial y}{\partial C}$ $=\frac{\frac{\partial f}{\partial x}-\frac{\partial g}{\partial x}}{\frac{\partial f}{\partial x}\cdot\frac{\partial g}{\partial y}-\frac{\partial f}{\partial y}\cdot\frac{\partial g}{\partial x}}>0$
which gives the second part as well.
6 Determining filters biases - demonstration
In this section we show an experimental example for the behavior of the functions $N(L,H)$, $P(L,H)$ and $\tilde{Sig}(L,H)$.
For the signal in the experiment, we used an incandescent lamp powered by the 50 Hz sinusoidal electrical grid, which gives 100 Hz periodic signal of the form $|\sin(x)|$. Then, we measured the signal by putting a EVK4-HD Prophesee’s event camera a few meters from the lamp. In order to simulate a small signal, we covered the lamp and made a small hole in the cover for the light. This way we got a signal of up to a few dozens of events per period. The background in the lab was kept dark during the experiment.
Then, we sampled the plane spanned by $L=$ bias_fo and $H=$ bias_hpf in steps of 10 units. For each sample we used an automatic feedback control algorithm that balanced the positive and negative event rate to be close to 50k events/sec and 100k events/sec in total. We remark that due to Theorem 1, the convergence of the algorithm to the desired background event rate was guaranteed. Conceptually, the algorithm performs a similar idea as the AGC feature in frame based cameras.
In Figure 6.1 one can see the value of bias_diff_on and bias_diff_off for each of the samples. It can be seen that in general, their value decreases when bias_hpf is higher and bias_fo is lower. In other words, one can see that when the filters are more aggressive, the noise is lower, and hence the algorithm can squeeze the sensitivity of the camera by lowering the values of the sensitivity biases, and still keeping a constant background event rate.
For each sample, after tuning the sensitivity biases, we recorded the signal for 0.5 second, getting 50 periods of the signal. Then, we counted the amount of positive and negative events in radius of 5 pixels around the center of the signal, and divided it by 50. In other words we defined $\tilde{Sig}(L,H)$ one time to be the amount of positive events per period around the center of the periodic signal given by the “covered” incandescent lamp, and one time to be the amount of negative events per period.
[…]
Figure 6.1: The values of the sensitivity biases $N=$ bias_diff_off and $P=$ bias_diff_on as function of the filters biases $L=$ bias_fo and $H=$ bias_hpf. In each sample of the filters biases we balanced and tuned the sensitivity biases to keep the BER close to the constant value of 100k events/sec.
Figure 6.2: The amount of positive and negative events per period of the signal originated by the incandescent lamp powered by the periodic electrical grid, as function of the filters biases.
Remark 4.
The tuning algorithm looped over the biases in straps by holding bias_fo, and tuning bias_hpf in steps of 10 units, from its lower value until reaching bias_hpf = 150. Hence, there was a big difference between the BER at the last sample in each strip, and the BER of the first sample in the next strip before the tuning. This difference can cause our tuning algorithm to converge unproperly in the beginning of each strip, and explains the valleys in the lower side of bias_hpf.
7 Conclusions
Solving the biases multivariable problem is a crucial stage one needs to pass to exploit the potential of event camera for performing any task. Hence, developing systematic heuristics and approaches to tackle this problem is an important challenge of the research community.
In this paper we presented a systematic heuristic to characterize the biases of an event camera for detection of small objects tasks in staring scenarios. We translated the biases behavior in this context into mathematical terms and showed that under suitable and reasonable assumptions, for any value of the filters biases, there exists exactly one pair of sensitivity biases that provide a predefined positive and negative event rate.
Having this understanding in mind, we demonstrated, in a certain scenario, under appropriate constraints on the event rate, how the sensitivity biases change, and how the signal of an incandescent lamp powered by the electrical grid is being influenced as well. This demonstraion gave us a better understanding on what can be done with event camera when one wants to detect signals with similar behavior.